OpenStudy (anonymous):

An ice sculpture is melting from the heat. Assuming its height changes by negative 5 over 32 meters every hour, what will be the change in height after 4 hours? negative 20 over 8 meters 5 over 8 meters negative 5 over 32 meters negative 5 over 8 meters

OpenStudy (jhannybean):

\[\large \sf -\frac{\dfrac{5}{32} ~m}{1~ hr}= \frac{x~m}{4~hr}\] cross multiply to find x meters

OpenStudy (anonymous):

Thanks! can you help me with this question

OpenStudy (anonymous):

Why is 3 + (−5) equal to −2? Because it is 5 units to the left of 0 on a horizontal number line Because it is 5 units to the right of 3 on a horizontal number line Because it is 5 units to the left of 3 on a horizontal number line Because it is 5 units to the right of 0 on a horizontal number line

OpenStudy (jhannybean):



OpenStudy (jhannybean):

I don't even know why that's a question, it seems ridicuous to me.

OpenStudy (jhannybean):

I'd just go with A.

OpenStudy (jhannybean):

Thought I could explain it but it's weird.